The net electric flux through a cubic box with sides that are 25.0 cm long is 4.70×103 N⋅m2/C . What charge ? is enclosed by the box?
Answer #1

Using Gauss Law

\text{Net Electric Flux}=\frac{Q_{enc}}{\epsilon_0}

or,

\\ Q_{enc}=\phi \epsilon_0 \\ Q_{enc}=4.7\times 10^3 (N.m^2/C)8.85\times 10^{-12} (C^2/N.m^2) \\ Q_{enc}=4.16\times 10^{-8} \ C